OpenStudy (joshmorris):

Marla bought one necklace and two bracelets for $67.50. Use the following information to determine the amount she paid for each bracelet: • The price of the necklace was $27. • The bracelets were both the same price.

OpenStudy (bonnieisflash1.0):

stop that

OpenStudy (bonnieisflash1.0):

stop changing your pic phonie

OpenStudy (bonnieisflash1.0):

you don't really have abs

OpenStudy (joshmorris):

Yes

OpenStudy (joshmorris):

Someone help

OpenStudy (shansha1):

1n + 2b = 67.50 n = 27 now we sub 27 in for n 27 + 2b = 67.50 2b = 67.50 - 27 2b = 40.50 b = 40.50/2 b = 20.25 check.. 1n + 2b = 67.50 1(27) + 2(20.25) = 67.50 27 + 40.50 = 67.50 67.50 = 67.50 (correct) so the bracelets (b) cost 20.25 each

OpenStudy (joshmorris):

Thanks

OpenStudy (shansha1):

welcome

OpenStudy (bonnieisflash1.0):

stop lying